# A Spectral Element Enrichment Wall-Model

### v3

@@ 1 Introduction
High Reynolds number turbulent flows are a norm, not the exception in engineering systems. Correctly simulating wall-bounded turbulent flows is crucial for applications, such as turbomachinery and propulsion/power engines [1]. In such applications, accurately resolving the velocity and thermal boundary layers at the wall is critical to compute the quantities of interest. However, prohibitive costs render computational fluid dynamics (CFD) simulations of these full-scale systems infeasible because they need to resolve smallest turbulent eddies.
@@
For most industry-relevant flow problems, this computational cost issue is addressed by using Reynolds-Averaged Navier-Stokes (RANS) simulations wherein all turbulent length and time scales are modeled and the mean flow solution is computed. For some problems, these methods are sufficient, but for unsteady simulations, simulations with flow separation, and flows with laminar-to-turbulent transition, RANS methods are unreliable. For such problems, the turbulence scales must be resolved throughout the domain in direct numerical simulations (DNS), or in the near-wall regions and modeled elsewhere in wall-resolved large-eddy simulations (WRLES). However, both of these approaches are computationally infeasible because the number of grid points in the near-wall region scales with $Re_{\tau}^{2}$ [2]. Hence, there is a need to develop reliable methods to reduce the cost of such simulations by modeling the smaller turbulent scales in the near-wall regions, but resolving the large-scale transients via wall-modeled large eddy simulations (WMLES).
@@
There have been a number of different WMLES algorithms developed in the literature. Most have been developed for low-order finite-difference (FD) and finite-volume (FV) methods [3], but only recently there have been efforts to develop such methods in the context of higher-order methods such as the Spectral Element Method (SEM) [4], Spectral Difference Method (SDM) [5], and the discontinuous Galerkin (DG) method [6] [7] [8]. Most of these methods follow the tradition of low-order FV methods and use the solution outside of the near-wall region to fit an empirical wall function in the boundary layer in order to compute the wall shear stress. Then, this shear stress is applied to the flow as the wall boundary condition. While these methods work well away from the wall, the modified boundary condition leads to inaccuracies in the near-wall region.
@@
This study seeks to make use of the high-order polynomial solution of SEM to develop a different kind of wall-model better suited for high-order methods than traditional models. The wall-model builds upon prior work on the development of so-called enriched basis turbulence wall models in the context of DG [8] [13] and the continuous finite element (FEM) [14] [15] [16]. These methods leverage the polynomial solution and augment it by considering the empirical wall function as part of the solution representation. Doing so prevents spurious oscillations in the near-wall region due to insufficient polynomial resolution of the boundary layer and improves the solution representation by allowing the degrees of freedom to represent the flow fluctuations instead of the mean flow behavior. This work serves to develop the framework for enrichment wall-models in SEM so that SEM because of its low dissipation and dispersion errors for its cost [9] and its computational efficiency and scalability [10].
@@ 3 Spectral Element Method
The governing equations are discretized using the spectral element method (SEM) in the Nek5000 [17] solver. In the SEM, the governing equations are put into a weighted residual formulation by multiplying Eqs. (1)-(2) by test functions $\bm{v}$ and $q$, respectively, and integrating over the elements, so the goal is to find $(\bm{u},p)\in X_{b}^{P}(\Omega)\times Y^{P}(\Omega)$ such that
@@ 4 Enrichment Method
The enriched SEM begins by putting the governing equations into weighted residual form as shown in Eqs. 3-4. However, for the enriched SEM, the solution is $(\bm{u},p)\in X_{b,\psi}^{P}(\Omega)\times Y^{P}(\Omega)$, where $X_{\psi}\subset H^{1}(\Omega)$ such that $X_{\psi}^{P}=\{u\,|\,u=u_{p}+\psi,u_{p}\in X^{P}(\Omega_{\psi})\cap X^{P}(\Omega_{p}),\psi\in H^{1}(\Omega_{\psi})\cap H^{1}(\Omega_{p}),\psi(\bm{x})=0\,\forall\bm{x}\in\Omega_{p}\}$, where $\Omega_{\psi}$ and $\Omega_{p}$ are the enriched and unenriched parts of the domain respectively, such that $\Omega_{\psi}\cup\Omega_{p}=\Omega$ and $\Omega_{\psi}$ and $\Omega_{p}$ do not overlap. $X_{b,\psi}$ is the subset of $X_{\psi}$ such that Dirichlet boundary conditions are satisfied on the boundary $\partial\Omega$. As is done in the traditional SEM, the weighted residual form is integrated by parts to give Eqs. 5-6. From here, the enriched solution representation from Eq. 24 is substituted into the weighted residual form:
@@
Because the enrichment function is constant within each timestep, the change in the enriched solution is equal to the change in the polynomial component: $\Delta\bm{u}^{n}=\Delta\bm{u}^{n}_{p}$. Further, Lemma 4.1 proves that $\Delta u_{p}^{n}\in H^{1}(\Omega)$. Hence, the standard SEM scatter-gather operations ensure that the solution is continuous. [27]:
@@ 5 Enrichment Function
where $y^{+}=h_{w}u_{\tau}/\nu$ is the non-dimensional distance from the wall. The law-of-the-wall function is fit to the resolved LES region by solving for the friction velocity that makes the modeled velocity equal to the the targeted LES velocity at the matching point, $u^{*}_{\|,(i,j)}$. The target LES velocity can be either the instantaneous or a time-averaged velocity [3]. In this study, it was seen that time-averaged velocities over a flow through time tended to be more stable than instantaneous values, so they will be used going forward. Using the $u^{*}_{\|,(i,j)}$, $h_{w,(i,j)}$, and $f_{w}$, an implicit nonlinear equation is solved for the friction velocity at the node, $u_{\tau,(i,j)}$:
@@
In order to accurately compute the integrals in Eqs. (32)-(34) a quadrature must be tailored for the enrichment function. The enrichment wall-model is a continuous and smooth function with large gradients in only the wall-normal direction. Hence, overintegration in the wall-normal direction is used. From testing, we determined that overintegrating using $1.5(P+1)$ GL quadrature points in the $\eta$ direction and standard $P+1$ GLL quadrature points in the $\xi$ and $\zeta$ directions was sufficient to accurately resolve the integrals for a range of polynomial orders and Reynolds numbers.
@@ 6 Numerical Results
$u=$ $Re_{\tau}\left(\frac{1}{k}\log{(1+ky^{+})}+(C-\frac{1}{k}\log{k})(1-e^{-\frac{y^{+}}{11}}-\frac{y^{+}}{11}e^{-\frac{y^{+}}{3}}\right)$ (53)
$+\epsilon\beta\sin(\alpha x)\cos(\beta z),$
@@ 6.1 Grid Resolution Study
In the more underresolved cases, cusps in the rms data can be seen at the end of the first element. For example, for $Re_{\tau}=2000$, rms data cusps can be seen at $y^{+}=200$. This location corresponds with the end of the wall-adjacent element, which is the matching point location. This behavior is a result of the discontinuous gradients because only $C^{0}$ continuity is enforced at element boundaries. It is not specific to the enrichment method, but is commonly seen in under-resolved SEM simulations [34] [35] [36]. It is a sign of an under-resolved mesh [34], which is expected because the large wall-adjacent element motivates the need for the wall model. Figures 5 and 6 show that this can be remedied with increased resolution, like increasing $P$. Despite this inaccuracy in the gradients, good agreement in the mean velocity profile is seen, highlighting the strength of the enrichment wall-model.
For all three $Re_{\tau}$ values, increasing $P$ and $N$ improves the agreement with the DNS as expected. Table 1 shows that large elements are being used near the wall in all cases. For all three Reynolds numbers, resolutions with the first element around $y_{1}^{+}=200$ show good agreement with the mean but under-predicted rms quantities. However, for $y_{1}^{+}\leq 100$ both the mean and rms quantities show good agreement with the DNS data. Hence, refining $N$ leads to more improvement than refining $P$, especially for the turbulent fluctuations. The larger Reynolds number cases require more refinement than the lower because of the larger gradients in the boundary layer profile. Overall, the enrichment wall-model effectively represents the near-wall behavior of the turbulent flow with large elements for a range of Reynolds numbers.
@@ 6.3 Comparison with Shear Stress Wall-Model
In this section, the enrichment wall-model is compared with purely polynomial standard SEM and a shear stress wall-modeled SEM to show the benefits. All three methods use the same initial conditions and problem specifications from Sec. 6. The standard SEM uses no-slip walls and no added filtering, so the numerics match the enriched case exactly. At the coarse near-wall resolutions presented, using a filter for added dissipation in the SEM leads to large oscillations and a mismatch of the targeted turbulent profile. The shear stress wall-model is implemented as described earlier in [4] [37]. The shear stress wall-model uses no-penetration slip-wall boundary conditions that apply a modeled shear stress at the wall. The matching point is taken as the top of the first wall-adjacent element. The model uses the Reichardt law-of-the-wall as the equilibrium wall-stress model. The initial solution is allowed to develop for $8$ flow-through times before the shear stress wall-model is activated. Additionally, an explicit filter with weight $0.05$ and cutoff ratio $0.65$ are used to add dissipation and greatly decrease the amount of slip at the wall. This added dissipation serves as the subgrid scale model for this LES, as is often done in SEM [35] [38].

### v4

@@ 1 Introduction
High Reynolds number turbulent flows are a norm, not the exception in engineering systems. Correctly simulating wall-bounded turbulent flows is crucial for applications, such as turbomachinery and propulsion/power engines [1]. In such applications, accurately resolving the velocity and thermal boundary layers at the wall is critical to compute the quantities of interest. However, prohibitive costs render computational fluid dynamics (CFD) simulations of these full-scale systems infeasible because they need to resolve the smallest turbulent eddies.
@@
For most industry-relevant flow problems, this computational cost issue is primarily addressed by using Reynolds-Averaged Navier-Stokes (RANS) simulations wherein all turbulent length and time scales are modeled and the mean flow solution is computed. For some problems, these methods are sufficient, but for unsteady simulations, simulations with flow separation, and flows with laminar-to-turbulent transition, RANS methods are unreliable. For such problems, the turbulence scales must be resolved throughout the domain in direct numerical simulations (DNS) or in the near-wall regions and modeled elsewhere in wall-resolved large-eddy simulations (WRLES). However, both of these approaches are often computationally infeasible because the number of grid points in the near-wall region scales with $Re^{2.05}$ for DNS and $Re^{1.86}$ for WRLES [2]. Only recently have LES methods begun to see use for industrial applications [3, 4, 5, 6]. Hence, there is a need to develop reliable methods to reduce the cost of such simulations by modeling the smaller turbulent scales in the near-wall regions, but resolving the large-scale transients via wall-modeled large eddy simulations (WMLES).
@@
There have been a number of different WMLES algorithms developed in the literature. Most have been developed for low-order finite-difference (FD) and finite-volume (FV) methods, [7] but only recently there have been efforts to develop such methods in the context of higher-order methods such as the Spectral Element Method (SEM) [8], Spectral Difference Method (SDM) [9], and the discontinuous Galerkin (DG) method [10, 11, 12]. Most of these methods follow the tradition of low-order FV methods and use the solution outside of the near-wall region to fit an empirical wall function in the boundary layer in order to compute the wall shear stress. Then, this shear stress is applied to the flow as the wall boundary condition. While these methods work well away from the wall, the modified boundary condition leads to inaccuracies in the near-wall region.
@@
This study seeks to make use of the high-order polynomial solution of SEM to develop a different kind of wall-model better suited for high-order methods than traditional models. The wall-model builds upon prior work on the development of so-called enriched basis turbulence wall models in the context of DG [12, 13] and the continuous finite element (FEM) [14, 15, 16]. These methods leverage the polynomial solution and augment it by considering the empirical wall function as part of the solution representation. Doing so prevents spurious oscillations in the near-wall region due to insufficient polynomial resolution of the boundary layer and improves the solution representation by allowing the degrees of freedom to represent the flow fluctuations instead of the mean flow behavior. This work serves to develop the framework for enrichment wall-models in SEM so that SEM because of its low dissipation and dispersion errors for its cost [17] and its computational efficiency and scalability [18].
@@ 3 Spectral Element Method
The governing equations are discretized using the spectral element method (SEM) in the Nek5000 solver[19]. In the SEM, the governing equations are put into a weighted residual formulation by multiplying Eqs. (1)-(2) by test functions $\bm{v}$ and $q$, respectively, and integrating over the elements, so the goal is to find $(\bm{u},p)\in X_{b}^{P}(\Omega)\times Y^{P}(\Omega)$ such that
@@ 4 Enrichment Method
The enriched SEM begins by putting the governing equations into weighted residual form as shown in Eqs. (3)-(4). However, for the enriched SEM, the solution is $(\bm{u},p)\in X_{b,\psi}^{P}(\Omega)\times Y^{P}(\Omega)$, where $X_{\psi}\subset H^{1}(\Omega)$ such that $X_{\psi}^{P}=\{u\,|\,u=u_{p}+\psi,u_{p}\in X^{P}(\Omega_{\psi})\cap X^{P}(\Omega_{p}),\psi\in H^{1}(\Omega_{\psi})\cap H^{1}(\Omega_{p}),\psi(\bm{x})=0\,\forall\bm{x}\in\Omega_{p}\}$, where $\Omega_{\psi}$ and $\Omega_{p}$ are the enriched and unenriched parts of the domain respectively, such that $\Omega_{\psi}\cup\Omega_{p}=\Omega$ and $\Omega_{\psi}$ and $\Omega_{p}$ do not overlap. $X_{b,\psi}$ is the subset of $X_{\psi}$ such that Dirichlet boundary conditions are satisfied on the boundary $\partial\Omega$. As is done in the traditional SEM, the weighted residual form is integrated by parts to give Eqs. (5)-(6). From here, the enriched solution representation from Eq. (24) is substituted into the weighted residual form:
@@
Because the enrichment function is constant within each timestep, the change in the enriched solution is equal to the change in the polynomial component: $\Delta\bm{u}^{n}=\Delta\bm{u}^{n}_{p}$. Further, Lemma 4.1 proves that $\Delta u_{p}^{n}\in H^{1}(\Omega)$. Hence, the standard SEM scatter-gather operations ensure that the solution is continuous [25]:
@@ 5 Enrichment Function
where $y^{+}=h_{w}u_{\tau}/\nu$ is the non-dimensional distance from the wall. The Spalding law-of-the-wall was also tested, but showed little effect on the solution and is computationally more expensive to evaluate [29]. The law-of-the-wall function is fit to the resolved LES region by solving for the friction velocity that makes the modeled velocity equal to the the targeted LES velocity at the matching point, $u^{*}_{\|,(i,j)}$. The target LES velocity can be either the instantaneous or a time-averaged velocity [7]. In this study, it was seen that time-averaged velocities over a flow through time tended to be more stable than instantaneous values, so they will be used going forward. Using the $u^{*}_{\|,(i,j)}$, $h_{w,(i,j)}$, and $f_{w}$, an implicit nonlinear equation is solved for the friction velocity at the node, $u_{\tau,(i,j)}$:
@@
In order to accurately compute the integrals in Eqs. (32)-(34) a quadrature must be tailored for the enrichment function. The enrichment wall-model is a continuous and smooth function with large gradients in only the wall-normal direction. Hence, overintegration in the wall-normal direction is used. From testing, we determined that overintegrating using $1.5(P+1)$ GL quadrature points in the $\eta$ direction and standard $P+1$ GLL quadrature points in the $\xi$ and $\zeta$ directions was sufficient to accurately resolve the integrals for a range of polynomial orders and Reynolds numbers. If the wall-model enrichment function or its derivative were discontinuous, a quadrature rule that could accurately integrate around the discontinuity, like a segmented quadrature rule, would be needed [13].
@@ 6 Numerical Results
$u=$ $Re_{\tau}\left(\frac{1}{k}\log{(1+ky^{+})}+(C-\frac{1}{k}\log{k})(1-e^{-\frac{y^{+}}{11}}-\frac{y^{+}}{11}e^{-\frac{y^{+}}{3}}\right)+\epsilon\beta\sin(\alpha x)\cos(\beta z),$ (53)
@@ 6.1 Grid Resolution Study
In the more underresolved cases, cusps in the rms data can be seen at the end of the first element. For example, for $Re_{\tau}=2000$, $N=20$, rms data cusps can be seen at $y^{+}=200$. This location corresponds with the end of the wall-adjacent element, which is the matching point location. The cusps are a result of the discontinuous gradients between the elements because only $C^{0}$ continuity is enforced at element boundaries. This behavior is not specific to the enrichment method, but is commonly seen in under-resolved high-order simulations [32, 33, 34]. It is a sign of an under-resolved mesh [32], which is expected because the large wall-adjacent element motivates the need for the wall model. Figures 5 and 6 show that the cusps can be remedied with increased resolution, like increasing $P$. Despite this inaccuracy in the gradients, good agreement in the mean velocity profile is seen, highlighting the strength of the enrichment wall-model.
For all three $Re_{\tau}$ values, increasing $P$ and $N$ improves the agreement with the DNS as expected. Table 1 shows that large elements are being used near the wall in all cases. For all three Reynolds numbers, resolutions with the first element around $y_{1}^{+}=200$ show good agreement with the mean but under-predicted rms quantities. However, for $y_{1}^{+}\leq 100$ both the mean and rms quantities show good agreement with the DNS data. If $N$ was refined such that $y_{1}^{+}<10$, the wall-model would have diminishing returns because the high-order polynomial solution will be more effective at capturing the near-wall region that does not include the log-layer. Larger values of $y_{1}^{+}$ lead to log-layer mismatch, as seen in the cases with $Re_{\tau}=2000$ and $N=10$. Hence, refining $N$ leads to more improvement than refining $P$, especially for the turbulent fluctuations. However, for a fixed mesh, increasing $P$ from $7$ to $9$ shows improvement in the RMS quantities, because the high-order polynomial can better represent the fluctuations near the wall. The larger Reynolds number cases require more refinement than the lower because of the larger gradients in the boundary layer profile. Overall, the enrichment wall-model effectively represents the near-wall behavior of the turbulent flow with large elements for a range of Reynolds numbers.
@@ 6.3 Comparison with Shear Stress Wall-Model
In this section, the enrichment wall-model is compared with purely polynomial standard SEM and a shear stress wall-modeled SEM to show the benefits. All three methods use the same initial conditions and problem specifications from Sec. 6. The standard SEM uses no-slip walls and no added filtering, so the numerics match the enriched case exactly. At the coarse near-wall resolutions presented, using a filter for added dissipation in the SEM leads to large oscillations and a mismatch of the targeted turbulent profile. The shear stress wall-model is implemented as described earlier in [8, 35]. The shear stress wall-model uses no-penetration slip-wall boundary conditions that apply a modeled shear stress at the wall. The matching point is taken as the top of the first wall-adjacent element. The model uses the Reichardt law-of-the-wall as the equilibrium wall-stress model. The initial solution is allowed to develop for $8$ flow-through times before the shear stress wall-model is activated. Additionally, an explicit filter with weight $0.05$ and cutoff ratio $0.65$ are used to add dissipation and greatly decrease the amount of slip at the wall. This added dissipation serves as the subgrid scale model for this LES, as is often done in SEM [33, 36].
@@ 7 Conclusions
Future work will develop the method further for more challenging flows such as higher Reynolds numbers and flows with separations. For higher Reynolds number flows where $y_{1}^{+}<200$ is not possible, we plan to develop a dissipative filter compatible with the enrichment method to address the log-layer mismatch and under-prediction of rms quantities that come from buildup of energy in the large wavenumbers. For simulations with flow separation, more complex wall-functions, such as differential equation based wall-functions, can be directly incorporated into the enrichment framework. Additionally, a sensor will be developed to disable the enrichment in regions of flow separation because the polynomial solution and no-slip boundary conditions should be sufficient to resolve the flow.